ROTATIONAL SPECTRUM AND STRUCTURE OF VAN DER WAALS COMPLEXES: CH_{2}F_{2}⋯CF_{2}=CH_{2} AND CH_{2}F_{2}⋯CF_{2}=CHF

YOSHIO TATAMITANI, YASUFUMI SUZUKI, \underline{TERUHIKO OGATA}, Department of Chemistry, Faculty of Science, Shizuoka University, Shizuoka, Japan 422-8529.

The microwave spectra of van der Waals complexes of difluoromethane(DFM) and 1,1-difluoroethene(DFE)/ trifluoroethene(TFE) CH_{2}F_{2}⋯CF_{2}=CH_{2} and CH_{2}F_{2}⋯CF_{2}=CHF have been observed using a Fourie transform microwave spectrometer from 5 to 19 GHz. The samples of 1% DFM and DFE/TFE in Ar were used to form the complexes in a super sonic expansion. For DFM⋯DFE, 48 a-type transitions were assigned and 72 a- and b-type transitions were assigned for DFM⋯TFE. The rotational and centrifugal distortion constants have been determined, leading to the molecular structures of DFM⋯DFE and DFM⋯TFE complexes. Both complexes were found to be C_{s} symmetry and bound with triple hydrogen bonds: One is the strong F⋯HC type and two are the weaker CH⋯F type. Molecular structures of these complexes are, however, quite different from each other. For DFM⋯DFE, the out of plane atoms are two hydrogens which are bonded to a fluorine in DFE and a fluorine in DFM is bonded to a hydrogen in DFE. For DFM⋯TFE, the out of plane atoms are two fluorines which attach to a hydrogen in DFE and a fluorine in DFM is bonded to the hydrogen in the same carbon of DFE. Ab initio MO calculations of MP2/6-311+G(d,p) lebel have supported these experimental results.